OpenStudy (anonymous):

An online company called “Animals Dressed as Humans” is making matching owner/dog socks for the upcoming winter. A total of 19 humans and dogs purchased socks. They sold a total of 60 individual socks. How many humans and dogs purchased socks to keep their feet warm? Set up the problem, explain why you set it up that way, and provide the solution.

OpenStudy (anonymous):

Allow x = humans and y = dogs. Then, x + y = 19 represents the total number of orders that were placed. 2x + 4y = 60 represents the total number of socks that were sold (2 because humans have two feet and 4 because dogs have four feet). Then, solve the first equation for x, to get x = 19 - y, and substitue this into the second equation for x. You get 2(19 - y) + 4y = 38 - 2y + 4y = 38 + 2y = 60. Subtract 38 from both sides to get 2y = 22 and divide both sides by 2 to get y = 11. Finally substitute this value back into the first equation for y, to get x + 11 = 19, and by subtracting 11 from both sides, x = 8. Therefore, 8 humans and 11 dogs purchased socks to keep their feet warm. To check, multiply 8(2) to get 16 and 11(4) to get 44, and 16 + 44 = 60?

OpenStudy (anonymous):

am i right?